periods​) After examining the various personal loan rates available to​ you, you find that you can borrow funds from a finance company at an APR of 9 percent compounded quarterly or from a bank at an APR of 10 percent compounded monthly. Which alternative is more​ attractive?
Answer #1

Ans We would prefer APR of 9 percent compounded quarterly since Effective Annual rate in this is less.

EAR = ( 1 + r )^n - 1
Compounded Quarterly
EAR= ( 1 + 9%/4)^4 - 1
EAR= 9.31%
Compounded Monthly
EAR= ( 1 + 10%/12)^12 - 1
EAR= 10.47%